OpenStudy (anonymous):

What is the simplified form of 9 times y to the seventh power over 15 times x to the fifth power * 5 x squared over 45 y cubed ?

OpenStudy (anonymous):

\[\frac{ 9y ^{7 }}{ 15x ^{5} }\times \frac{ 5x ^{2} }{ 45y ^{3} }\]

OpenStudy (anonymous):

If this were a set of fractions without variables, would you be able to solve it?

OpenStudy (anonymous):

i just need the answer but yeah probably

OpenStudy (anonymous):

I don't give answers. Sorry.

OpenStudy (jdoe0001):

\(\bf \cfrac{ 9y ^{7 }}{ 15x ^{5} }\times \cfrac{ 5x ^{2} }{ 45y ^{3} } \implies \cfrac{ 9yyyyyyy}{ (3)(5)xxxxx }\times \cfrac{ 5xxx }{ (9)(5)yyy }\) cancel out like-terms